Please help me: a candy company is going to test market 1 kg packages each containing 3 kinds of candy the company plans to prepare 10,000 packages each will contain mints that sell for $2.20 per kg caramels that sell for $3.00 per kg and toffee that sells for $3.20 per kg marketing says that they must use twice as many caramels as toffee in each package the cost will be $2.72 per kg to break even how many total kg of mints will they need for the 10,000 packages
Answer by scott8148(5891)

let x="kg of toffee", so 2x="kg of caramels" and 1-3x="kg of mints"

x(3.20) + 2x(3.00) + (1-3x)(2.20) = 2.72

expanding __ 3.2x + 6x + 2.2 - 6.6x = 2.72

subtracting 2.2 and combining terms __ 2.6x = .52

dividing by 2.6 __ x = .2

substituting __ 1-3(.2) = .4 __ this is the kg of mints in one package

so they will need .4(10000) or 4000kg of mints in total